# Pose-Diversified Augmentation with Diffusion Model for Person Re-Identification

### v1

@@ Abstract
Person re-identification (Re-ID) often faces challenges due to variations in human poses and camera viewpoints, which significantly affect the appearance of individuals across images. Existing datasets frequently lack diversity and scalability in these aspects, hindering the generalization of Re-ID models to new camera systems. Previous methods have attempted to address these issues through data augmentation; however, they rely on human poses already present in the training dataset, failing to effectively reduce the human pose bias in the dataset. We propose Diff-ID, a novel data augmentation approach that incorporates sparse and underrepresented human pose and camera viewpoint examples into the training data, addressing the limited diversity in the original training data distribution. Our objective is to augment a training dataset that enables existing Re-ID models to learn features unbiased by human pose and camera viewpoint variations. To achieve this, we leverage the knowledge of pre-trained large-scale diffusion models. Using the SMPL model, we simultaneously capture both the desired human poses and camera viewpoints, enabling realistic human rendering. The depth information provided by the SMPL model indirectly conveys the camera viewpoints. By conditioning the diffusion model on both the human pose and camera viewpoint concurrently through the SMPL model, we generate realistic images with diverse human poses and camera viewpoints. Qualitative results demonstrate the effectiveness of our method in addressing human pose bias and enhancing the generalizability of Re-ID models compared to other data augmentation-based Re-ID approaches. The performance gains achieved by training Re-ID models on our offline augmented dataset highlight the potential of our proposed framework in improving the scalability and generalizability of person Re-ID models. The project page is available at https://ku-cvlab.github.io/Diff-ID/.
Figure 1: Teaser: Generated samples from a reference identity by Diff-ID. Our method can generate new human images conditioned on various poses while maintaining the appearance of the original human identity, which are used for augmenting the training data for person re-identification (Re-ID) task. Ref. denotes the reference image.
1 Introduction
@@
Person re-identification (Re-ID) has been popularly used to retrieve [67] and track [53] individuals across images captured by multiple cameras with non-overlapping views [6, 66]. It has applications in various fields, such as robotics [8] and surveillance [53].
@@
However, real-world deployment of Re-ID often faces several challenges. A major obstacle arises from the differences in human pose [7, 45, 63] and camera viewpoint [1, 27] within images of the same individual. These variations significantly affect the appearance of individuals, making the identification task difficult. To overcome this, existing models often require diverse images, but datasets for Re-ID are limited in terms of scalability and generalizability [10], since manual labeling of individuals across cameras is notoriously challenging [34, 64]. Furthermore, the limited number of cameras in the datasets makes it challenging to generalize models to new camera networks [37, 53, 72]. As a result, learning pose-invariant features for Re-ID that remain consistent across different human poses and camera viewpoints remains a challenge.
@@
To address these challenges, previous attempts [9, 10, 13, 24, 33, 38, 69, 70, 73] have involved data augmentation aimed at increasing the diversity of training data. Early methods employed simple image transformation techniques [24, 38, 73], while more recent approaches leverage generative adversarial networks (GAN)-based models for augmentation [33, 69, 70]. However, since these methods rely on poses already existing in the Re-ID dataset to generate augmented images, they do not effectively address the core issue of pose bias in the training data distribution. Even in the case that the human pose is the same, differences in camera angles may induce significant changes in appearance [57]. For example, a picture taken from a high-angle camera viewpoint may appear different from one taken from a low-angle viewpoint, potentially causing the same person to be identified as different individuals. To generalize and reduce the training data bias, it is crucial to diversify the training data’s human pose and camera viewpoint distribution.
In this paper, we define a novel pose augmentation techniques to consider both human pose and camera viewpoint. We propose a diffusion model-based framework, called Diff-ID, to augment the training data with poses and camera viewpoints that are underrepresented, i.e., rarely appeared, in the original training dataset. By including these less common examples, we aim to minimize the impact of pose and viewpoint variations on appearance. This approach focuses to generalize the distribution of the training dataset, enabling the training of more scalable and generalizable Re-ID models. Furthermore, our model architecture is designed to consider both human pose and camera viewpoint, acknowledging their joint influence on a person’s appearance in Re-ID.
@@
However, generating new images from sparse and limited poses is a challenging task. Since generative models learn the training data distribution, they may struggle to generate images with infrequently seen poses in the training data distribution, even if they are well-trained on a Re-ID dataset. To address this issue, we leverage Stable Diffusion [42], a pre-trained large-scale generative model, to augment our data with images of sparse and limited poses. By effectively leveraging the pre-trained knowledge, we can successfully augment the data, even when the given human pose and camera condition are uncommon in the training dataset. Also, to jointly provide human pose and camera viewpoint that are rarely seen in the training dataset, we use the SMPL [35] model. We conditioned the generative model with SMPL, enabling realistic human rendering and providing implicit information about the camera viewpoint through depth data.
@@
Our objective is not developing a new Re-ID model, but rather augmenting a training dataset that enables Re-ID models to learn features unbiased to pose variations. We present performance gains achieved by training the Re-ID model on our offline augmented dataset using our proposed architecture. Furthermore, qualitative results demonstrate that, compared to other approaches, our architecture generates significantly more realistic and feasible images for human poses and viewpoints that were not well-represented in existing Re-ID models. This highlights the effectiveness of our method in addressing the pose bias issue and enhancing the generalizability of Re-ID models.
2 Related Work
@@
Figure 2: Intuition of our framework. Upon observing the highly biased viewpoint and human pose distributions in training dataset, we augment the dataset by manipulating SMPL body shapes and feeding the rendered shapes into a generative model to fill in sparsely distributed poses and viewpoints. With this augmented dataset, we can train a Re-ID model that is robust to viewpoint and human pose biases.
@@
Person Re-ID is a crucial task in computer vision that involves identifying individuals across different camera views. Recent advancements in Re-ID [5, 10, 14, 19, 26, 31, 46, 51, 54] leverage deep learning techniques to improve Re-ID performance. The most significant method is Distance Metric Learning [20, 30, 32, 55, 58, 68], such as Triplet Loss [22], which utilizes deep learning models to project images into a feature embedding space where the vectors of the same identity are close together, while those of different identities are far apart.
@@
Various approaches [9, 38, 24, 39, 70, 33, 69, 73] employed data augmentation techniques to enhance person Re-ID models. Operations such as random resizing, cropping, and horizontal flipping [38], along with adversarially occluded samples [24], increase the diversity of training data. Approaches using methods such as DropBlock [9] and background transformation [39] show further improved feature learning. Meanwhile, there are approaches [70, 33, 69] using a generative model, which is Generative Adversarial Networks (GANs) [15], for data augmentation in Re-ID task. DCGANs [70] generate new samples, enriching training datasets which improves the performance of the Re-ID model.  [33] utilizes a data augmentation method involving pose-transferred sample augmentations. In  [69], data augmentation is achieved through a joint learning framework that couples generative and discriminative modules. Techniques like random erasing [73] can also improve model robustness and generalization.
@@
Pose-guided person image generation task aims to generate a person’s image with a specific pose. With the promising result of diffusion model [21, 48], there have been approaches [2, 61, 42, 17, 23, 28, 56, 75] that aim to generate human images with desired structure by applying diffusion model with 2D human pose as a condition. PIDM [2] is the first method that integrates the diffusion model [21, 48] pose-guided person synthesis task, using OpenPose [3] as the pose condition. ControlNet [61] takes a method to fine-tune Stable Diffusion [42] using zero-initialized convolution layer to inject 2D human pose condition. Addressing the challenge caused by sparse pose conditions, PoCoLD [17] adopts DensePose [16] as a robust pose representation. Meanwhile, pose-guided person video generation is also noticeable. Recent works such as DreamPose [28] and AnimateAnyone [23] generate a video guided by pose-sequence which is temporally coherent, and use the fine-tuning method. MagicAnimate [56] also generates a video guided by pose-sequence conditions, but leverages pose ControlNet [61] as a pose injector. Champ [75], a subsequent work of AnimateAnyone [23] and MagicAnimate [56], uses SMPL [36] guidance instead of only 2D human pose.
3 Method
Figure 3: Visualization of the effect of viewpoint and human pose augmentation. We compare visualizations of camera viewpoint and human pose distributions for the Market-1501 [65] and DukeMTMC-reID datasets [71]. The left figures (i) display the camera viewpoint distribution derived from SMPL [35], while the right figures (ii) illustrate the pose distribution. In (i), from left to right, we show the viewpoint distributions of the training dataset, the augmented dataset, and the combination of both. Similarly, in (ii), from left to right, we present t-SNE [50] visualizations of the human pose distributions, showing poses from the training dataset, poses used for augmentation, and the combination of both. These visualizations demonstrate that our pose augmentation successfully diversifies both viewpoint and human pose distributions.
3.1 Motivation and Overview
@@
In this paper, we tackle the challenging Re-ID problem of limited camera viewpoints and human poses in training datasets. Re-ID datasets typically contain a limited number of cameras, ranging from 6 in Market-1501 [65] to 8 in the DukeMTMC-reID dataset [71]. This small number of fixed cameras capture individuals within a restricted range of viewpoints, significantly impacting the generalization performance of Re-ID models. Furthermore, since the cameras mostly capture people walking or standing still, this also affects the generalization capabilities of Re-ID models. The analysis in Figure 2 illustrates the restricted camera viewpoints and human poses captured in the training dataset.
@@
We aim to diversify the range of camera viewpoints and human poses by augmenting the training dataset with generative models. However, generative models sample from the training dataset distribution. Without explicit consideration for generalization, the quality of generated images with camera viewpoints and human poses that are rarely present in the training dataset distribution may be compromised. To address this, we leverage the powerful generalization capability of a foundation generative model trained on a billion-scale dataset, i.e, Stable Diffusion [42]. We carefully select and integrate components to enable camera viewpoint and human pose conditioning within the generative model for augmentation. By leveraging its generalization ability and applying viewpoint conditioning, we can effectively augment the training dataset with diverse camera viewpoints and human poses, as shown in Figure 3.
@@
In this section, we explain the augmentation process in Sec. 3.2. Subsequently, we present human pose and camera viewpoint condition from SMPL in Sec. 3.3, and how to inject condition with identity into diffusion model in Sec. 3.4.
3.2 Pose-Diversified Augmentation with Diffusion Models
To mitigate the biased camera viewpoint and human pose in training dataset, we augment images with uniformly distributed camera viewpoint and human poses sourced from outside the training dataset. Specifically, for camera viewpoints, we augment the images adjusting two factor: elevation and azimuth. For the elevation, we uniformly sample from $0^{\circ}$ to $30^{\circ}$, assuming that the camera is not positioned below the ground and that a person becomes indistinguishable when the camera is positioned above a certain degree. For the azimuth, we randomly sample from $0^{\circ}$ to $360^{\circ}$, as a person can be captured in any random direction. With these consideration, we found the distribution of camera viewpoint is significantly unbiased, as depicted in Figure 2. For human poses, we source them from outside the training dataset for diversification, e.g., using poses extracted from dance videos [4]. This significantly diversifies the human poses, facilitating generalization to various poses during testing. Given predefined camera and pose distributions sourced from outside the training dataset, we aim to augment the training dataset using a generative model conditioned on the camera and pose.
Figure 4: Overall architecture of Diff-ID. Given the viewpoint and pose distributions, we first render the body shape sampled from the distribution using SMPL [36], generating the corresponding skeleton, depth map, and normal maps. These conditions, along with a reference image for identity preservation, are then fed into Diff-ID, which consists of two branches: the reference U-Net processes the identity information from the reference image, while the denoising U-Net generates a person with the same identity, given the input conditions. The denoising U-Net generates images by iterating through the denoising process.
3.3 Human Pose and Camera Viewpoint Condition from SMPL
@@
To augment the data with varying human pose and camera viewpoint, we provide specific conditions to the diffusion model. Previous works [11, 40, 49] have primarily addressed pose control by providing a human pose skeleton. However, relying solely on the human skeleton has limitations due to missing information: When projected onto 2D images, the human skeleton lacks depth information, posing ambiguity for the model when inferring viewpoint information from the skeleton. For example, if the camera is above a person, the skeleton would appear compressed vertically. Without depth information, the model cannot distinguish whether the camera is positioned above the person or the person is simply short.
@@
In this regard, in addition to the human pose skeleton, we utilize SMPL [36], a human body model used for realistic human rendering. SMPL [36] can model intricate human shapes, including complex body articulations in 3D space. From this human model, we can easily extract 2D representations of the 3D human by rendering the model, such as depth maps which implicitly contain camera viewpoint information.
To provide rich information about camera viewpoint, we use depth maps from SMPL [36] as a condition for the diffusion model, along with the human skeleton and a photo to control identity. Additionally, we incorporate surface normals from SMPL [36], as they encode detailed human surface information and can help generate more precise augmentations. Gathering these conditions, we feed them to diffusion model as a guidance.
The conditions can be classified into two categories: 1) Conditions that should be spatially aligned with the generated output, such as depth information from SMPL [36] and the human skeleton. 2) Conditions that are not necessarily spatially aligned with the generated output, such as identity information. While maintaining the original U-Net [43] structure to preserve its pre-trained knowledge, we carefully add components to the Stable Diffusion model to handle these two different types of conditions.
3.4 Conditional Diffusion Models for Pose-Diversified Augmentation

### v2

# Pose-dIVE: Pose-Diversified Augmentation for Person Re-Identification

@@ Abstract
Person re-identification (Re-ID) often faces challenges due to variations in human poses and camera viewpoints, which significantly affect the appearance of individuals across images. Existing datasets frequently lack diversity and scalability in these aspects, hindering the generalization of Re-ID models to new camera systems. We propose Pose-dIVE, a novel data augmentation approach that incorporates sparse and underrepresented human pose and camera viewpoint examples into the training data, addressing the limited diversity in the original training data distribution. Our objective is to augment the training dataset to enable existing Re-ID models to learn features unbiased by human pose and camera viewpoint variations. To achieve this, we leverage the knowledge of pre-trained large-scale diffusion models. By conditioning the diffusion model on both the human pose and camera viewpoint concurrently through the SMPL model, we generate training data with diverse human poses and camera viewpoints. Experimental results demonstrate the effectiveness of our method in addressing human pose bias and enhancing the generalizability of Re-ID models compared to other data augmentation-based Re-ID approaches.
footnotetext: Equal contribution.
Introduction
@@
Person re-identification (Re-ID) has been popularly used to retrieve (Zheng, Yang, and Tian 2017) and track (Wang 2013) individuals across images captured by multiple cameras with non-overlapping views (Chen et al. 2018; Zheng, Yang, and Hauptmann 2016). It has applications in various fields, such as robotics (Coşar and Bellotto 2020) and surveillance (Wang 2013).
@@
However, real-world deployment of Re-ID often faces several challenges. A major obstacle arises from the differences in human pose (Cho and Yoon 2016; Sarfraz et al. 2018; Zhao et al. 2017) and camera viewpoint (Bak et al. 2014; Karanam, Li, and Radke 2015) within images of the same individual. These variations significantly affect the appearance of individuals, making the identification task difficult. To overcome this, existing models often require diverse images, but datasets for Re-ID are limited in terms of scalability and generalizability (Fu et al. 2021), since manual labeling of individuals across cameras is notoriously challenging (Liu et al. 2014; Zhao, Ouyang, and Wang 2013). Furthermore, the limited number of cameras in the datasets makes it challenging to generalize models to new camera networks (Luo, Song, and Zhang 2020; Wang 2013; Zhong et al. 2018). As a result, learning pose-invariant features for Re-ID that remain consistent across different human poses and camera viewpoints remains a challenge.
Figure 1: Pose-dIVE diversifies the viewpoint and human pose of the Re-ID dataset to help generalize and improve the performance of arbitrary Re-ID models.
Figure 2: Visualization of the effect of viewpoint and human pose augmentation. We compare visualizations of camera viewpoint and human pose distributions for the Market-1501 (Zheng et al. 2015). The left figures (i) display the camera viewpoint distribution derived from SMPL, while the right figures (ii) illustrate the pose distribution. In (i), from left to right, we show the viewpoint distributions of the training dataset, the augmented dataset, and the combination of both. Similarly, in (ii), from left to right, we present t-SNE (Van der Maaten and Hinton 2008) visualizations of the human pose distributions, showing poses from the training dataset, followed by augmented poses sourced from outside the dataset. These visualizations demonstrate that our pose augmentation successfully diversifies both viewpoint and human pose distributions.
@@
To address these challenges, previous attempts (Dai et al. 2019; Fu et al. 2021; Gong, Huang, and Chen 2021a; Huang et al. 2018; Liu et al. 2018; Luo et al. 2019; Zheng et al. 2019) have involved data augmentation aimed at increasing the diversity of training data. Early methods employed simple image transformation techniques (Huang et al. 2018; Luo et al. 2019; Zhong et al. 2020), while more recent approaches leverage generative adversarial networks (GAN)-based models for augmentation (Liu et al. 2018; Zheng et al. 2019). However, since these methods rely on poses already existing in the Re-ID dataset to generate augmented images, they do not effectively address the core issue of pose bias in the training data distribution. Even in the case that the human pose is the same, differences in camera angles may induce significant changes in appearance (Ye et al. 2021). For example, a picture taken from a high-angle camera viewpoint may appear different from one taken from a low-angle viewpoint, potentially causing the same person to be identified as different individuals. To generalize and reduce the training data bias, it is crucial to diversify the training data’s human pose and camera viewpoint distribution.
In this paper, we define a novel pose augmentation technique which enables Re-ID models to robustly handle pose (i.e., human pose and camera viewpoint) variation, as illustrated in Figure 1. We propose a diffusion model-based framework, called Pose-dIVE, to augment the training dataset with generated images that contain human poses and camera viewpoints relatively underrepresented (i.e., rarely appearing) within the original training dataset. This effectively normalizes the distribution of the training dataset, allowing Re-ID models trained to be more generalizable and scalable.
@@
However, a generative model naively trained under a Re-ID dataset would struggle to generate such images with infrequent poses and cameras, as such a network would likewise reflect the bias within the training data distribution. To overcome this issue, we leverage Stable Diffusion (Rombach et al. 2022), a pre-trained large-scale generative model, whose powerful generative priors enable realistic, robust generation of images with human pose and camera conditions uncommon in the training dataset. Also, to jointly provide human pose and camera viewpoint that are rarely seen in the training dataset, we use the SMPL (Loper et al. 2015) model. We condition the generative model with SMPL, enabling realistic human rendering and providing implicit information about the camera viewpoint through depth data.
@@
Our method focuses on augmenting a training dataset in a manner that enables Re-ID models to learn features unbiased to pose variations. We present performance gains achieved by training the Re-ID model on our offline augmented dataset using our proposed architecture.
This highlights the effectiveness of our method in addressing the pose bias issue and enhancing the generalizability of Re-ID models.
@@
Figure 3: Pose-dIVE framework. Upon observing the highly biased viewpoint and human pose distributions in training dataset, we augment the dataset by manipulating SMPL body shapes and feeding the rendered shapes into a generative model to fill in sparsely distributed poses and viewpoints. With this augmented dataset, we can train a Re-ID model that is robust to viewpoint and human pose biases.
Related Work
@@
Person Re-ID is a crucial task in computer vision that involves identifying individuals across different camera views. Recent advancements in Re-ID (Chen et al. 2023; Fu et al. 2021; Gong, Huang, and Chen 2021b; He et al. 2021b; Kapil 2021; Li, Sun, and Li 2023; Somers, De Vleeschouwer, and Alahi 2023; Wang et al. 2019; Wieczorek, Rychalska, and Dabrowski 2021) leverage deep learning techniques to improve Re-ID performance. The most significant method is Distance Metric Learning (Hirzer et al. 2012; Koestinger et al. 2012; Liao and Li 2015; Xiong et al. 2014; Yu, Wu, and Zheng 2018; Zheng, Gong, and Xiang 2011), such as Triplet Loss (Hoffer and Ailon 2015), which utilizes deep learning models to project images into a feature embedding space where the vectors of the same identity are close together, while those of different identities are far apart.
@@
Various approaches (Dai et al. 2019; Luo et al. 2019; Huang et al. 2018) employed data augmentation techniques to enhance person Re-ID models. Operations such as random resizing, cropping, and horizontal flipping (Luo et al. 2019), along with adversarially occluded samples (Huang et al. 2018), increase the diversity of training data. Approaches using methods such as DropBlock (Dai et al. 2019) and background transformation (McLaughlin, Del Rincon, and Miller 2015) show further improved feature learning. Meanwhile, there are approaches (Zheng, Zheng, and Yang 2017; Liu et al. 2018; Zheng et al. 2019) using a generative model, which is Generative Adversarial Networks (GANs) (Goodfellow et al. 2020), for data augmentation in Re-ID task. DCGANs (Zheng, Zheng, and Yang 2017) generate new samples, enriching training datasets which improves the performance of the Re-ID model. Liu et al. (2018) utilizes a data augmentation method involving pose-transferred sample augmentations. In Zheng et al. (2019), data augmentation is achieved through a joint learning framework that couples generative and discriminative modules.
@@
Pose-guided person image generation task aims to generate a person’s image with a specific pose. With the promising result of diffusion model (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020), there have been approaches (Bhunia et al. 2023; Zhang, Rao, and Agrawala 2023; Rombach et al. 2022; Han et al. 2023; Hu et al. 2023; Karras et al. 2023; Xu et al. 2023; Zhu et al. 2024) that aim to generate human images with desired structure by applying diffusion model with 2D human pose as a condition. Leveraging the extensive pre-trained knowledge of Stable Diffusion (Rombach et al. 2022), these methods demonstrate impressive generalizability across various modalities and human poses. In this work, our goal is not to develop a superior pose-conditioned diffusion model, but rather to propose a framework that utilizes the inherent generalizability of diffusion models for the task of Re-ID.
Method
Motivation and Overview
@@
In this paper, we tackle the challenging Re-ID problem of limited camera viewpoints and human poses in training datasets. Re-ID datasets typically contain a limited number of cameras, ranging from 6 in the CUHK03 dataset (Li et al. 2014), 6 in Market-1501 (Zheng et al. 2015) to 15 in the MSMT17 dataset (Wei et al. 2018). Note that MSMT17 has the largest number of cameras among the publicly available Re-ID datasets. This small number of fixed cameras capture individuals within a restricted range of viewpoints, significantly impacting the generalization performance of Re-ID models. Furthermore, since the cameras mostly capture people with limited variations in human poses, such walking or standing still, this also affects the generalization capabilities of Re-ID models. The Figure 2 illustrates the restricted camera viewpoints and human poses captured in the training dataset.
These limitations in datasets have been a long-standing crucial problem in Re-ID models (Ye et al. 2021). Early research efforts attempted to solve this challenge of pose and camera variations through algorithmically or hand-crafted feature construction based on body structures (Zhong et al. 2018; Somers, De Vleeschouwer, and Alahi 2023). However, the restricted camera viewpoints and human poses are inherent issues in the dataset itself, and solely relying on algorithmic or hand-crafted methods has limitations in fully resolving these problems.
@@
We aim to diversify the range of camera viewpoints and human poses by augmenting the training dataset with generative models. However, as generative models sample from the training dataset distribution, without explicit consideration for generalization, the quality of generated images with camera viewpoints and human poses that are rarely present in the training dataset distribution may be compromised. To address this, we leverage the powerful generalization capability of a foundation generative model trained on a billion-scale dataset, i.e, Stable Diffusion (Rombach et al. 2022). We carefully select and integrate components to enable camera viewpoint and human pose conditioning within the generative model for augmentation. By leveraging its generalization ability and applying viewpoint conditioning, we can effectively augment the training dataset with diverse camera viewpoints and human poses, as shown in Figure 3.
@@
In this section, we first explain the augmentation process. Subsequently, we describe how human pose and camera viewpoint conditions are derived from the SMPL model, and how to incorporate these conditions along with identity information into the diffusion model.
Human Pose and Camera Viewpoint Condition with SMPL
@@
To augment the data with varying human pose and camera viewpoint, we provide specific conditions to the generative model. Previous works (Ge et al. 2018; Qian et al. 2018; Tang et al. 2020) have primarily addressed pose control by providing a human pose skeleton. However, relying solely on the human skeleton has limitations due to missing information: when projected onto 2D images, the human skeleton lacks depth information, posing ambiguity for the model when inferring viewpoint information from the skeleton. For example, if the camera is above a person, the skeleton would appear compressed vertically. Without depth information, the model cannot distinguish whether the camera is positioned above the person or the person is simply short.
@@
In this regard, in addition to the human pose skeleton, we utilize SMPL (Loper et al. 2023), a human body model used for realistic human rendering. SMPL can model intricate human shapes, including complex body articulations in 3D space. From this human model, we can easily extract 2D representations of the 3D human by rendering the model, such as depth maps which implicitly contain camera viewpoint information.
To enrich the camera viewpoint information, we incorporate depth maps. Additionally, surface normals from SMPL are used to capture detailed human surface characteristics, enhancing the precision of the generated augmentations. As a result, depth maps, surface normals, and human skeletons from SMPL, along with a reference image to control identity, are fed into the generative model as guidance.
Pose Diversified Augmentation
To mitigate the biased camera viewpoint and human pose in training dataset, we augment images with uniformly distributed camera viewpoint and human poses sourced from outside the training dataset. For camera viewpoints, we augment the images adjusting two factor: elevation and azimuth. For the elevation, we uniformly sample from $0^{\circ}$ to $30^{\circ}$, assuming that the camera is not positioned below the ground and that a person becomes indistinguishable when the camera is positioned above a certain degree. For the azimuth, we randomly sample from $0^{\circ}$ to $360^{\circ}$, as a person can be captured in any random direction. With these considerations, we found the distribution of camera viewpoint is significantly unbiased, as depicted in Figure 2. For human poses, we source them from outside the training dataset for diversification, e.g., using poses extracted from dance videos (Chan et al. 2019). As Re-ID datasets often contain similar human poses, solely training on these datasets can lead to overfitting and limited generalization capabilities. By incorporating these diverse external poses, our approach enhances the model’s ability to handle unseen poses, thereby addressing this limitation and improving overall performance. Using these diversified viewpoint and pose distributions, we render SMPL models into 2D representations to serve as conditions for the generative model.
Pose Diversified Generation with Stable Diffusion